You have just been hired as a financial analyst for Lydex Company, a manufacturer of safety helmets. Your boss has asked you to perform a comprehensive analysis of the company’s financial statements, including comparing Lydex’s performance to its major competitors. The company’s financial statements for the last two years are as follows:

Lydex Company
Comparative Balance Sheet
This Year Last Year
Assets
Current assets:
Cash $ 880,000 $ 1,120,000
Marketable securities 0 300,000
Accounts receivable, net 2,380,000 1,480,000
Inventory 3,520,000 2,200,000
Prepaid expenses 240,000 180,000
Total current assets 7,020,000 5,280,000
Plant and equipment, net 9,360,000 8,970,000
Total assets $ 16,380,000 $ 14,250,000
Liabilities and Stockholders' Equity
Liabilities:
Current liabilities $ 3,930,000 $ 2,820,000
Note payable, 10% 3,620,000 3,020,000
Total liabilities 7,550,000 5,840,000
Stockholders' equity:
Common stock, $75 par value 7,500,000 7,500,000
Retained earnings 1,330,000 910,000
Total stockholders' equity 8,830,000 8,410,000
Total liabilities and stockholders' equity $ 16,380,000 $ 14,250,000
Lydex Company
Comparative Income Statement and Reconciliation
This Year Last Year
Sales (all on account) $ 15,780,000 $ 12,780,000
Cost of goods sold 12,624,000 9,585,000
Gross margin 3,156,000 3,195,000
Selling and administrative expenses 1,794,000 1,572,000
Net operating income 1,362,000 1,623,000
Interest expense 362,000 302,000
Net income before taxes 1,000,000 1,321,000
Income taxes (30%) 300,000 396,300
Net income 700,000 924,700
Common dividends 280,000 462,350
Net income retained 420,000 462,350
Beginning retained earnings 910,000 447,650
Ending retained earnings $ 1,330,000 $ 910,000

To begin your assignment you gather the following financial data and ratios that are typical of companies in Lydex Company’s industry:

Current ratio 2.3
Acid-test ratio 1.0
Average collection period 30 days
Average sale period 60 days
Return on assets 8.4 %
Debt-to-equity ratio 0.7
Times interest earned ratio 5.7
Price-earnings ratio 10

2. You decide next to assess the company’s stock market performance. Assume that Lydex’s stock price at the end of this year is $78 per share and that at the end of last year it was $46. For both this year and last year, compute: (Round your "Percentage" answers to 1 decimal place and other intermediate and final answers to 2 decimal places.)

a. The earnings per share.

b. The dividend yield ratio.

c. The dividend payout ratio.

d. The price-earnings ratio.

e. The book value per share of common stock.
Answer #1

ANSWER:

(A)

Earning per share = net income/number of common stockholders

for this year,

Earning per share = $700000/100000 = $7 per share

where, number of common stockholders = $7500000/$75 = 100000

For last year,

Earning per share = $924700/100000 = $9.25 per share

where, number of common stockholders = $7500000/$75 = 100000

(B)

Dividend yield ratio = dividend per share/market price per share

for this year,

dividend yield ratio = ($280000/100000)/$78 = 3.6%

for last year,

dividend yield ratio = ($462350/100000)/$46 = 10.0%

(C)

Dividend payout ratio = dividend per share/earning per share

for this year,

dividend payout ratio = ($280000/100000)/$7 = 40%

for last year,

dividend payout ratio = ($462350/100000)/$9.25 = 49.9%

(D)

Price-earning ratio = market price per share/earning per share

for this year,

Price-earning ratio = $78/$7 = 11.14 times

for last year,

Price-earning ratio = $46/$9.25 = 4.97 times

(E)

Book value per share = total stockholders equity/number of common stock

for this year,

Book value per share = $8830000/100000 = $88.30 per share

for last year,

Book value per share = $8410000/100000 = $84.10 per share